OpenStudy (anonymous):

i need the formula for this word problem....a train travels 20 feet in 1/5 second. at this same speed, how many feet will it travel in three seconds.

OpenStudy (anonymous):

Its speed is \(\frac{20}{1/5}=100 ft/s\). So in three seconds, it will travel \(100(3)=300 ft\).

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

i have another one please

OpenStudy (anonymous):

Ask away!

OpenStudy (anonymous):

in printing an article of 48,000 words, a printer decides to use two sizes of type. using the larger type, a printed page contains 1800 words. using the smaller type, a page contains 2400 words. the article is allotted 21 full pages in a magazine. how many pages must be in smaller type? please give me the formula

OpenStudy (anonymous):

All right. You can use the givens in the problem to write two equations and then solve them to get the number of each type. Let x be the number of the smaller type and y the number of the bigger one. We know that the total number of pages should be 21. So, \(x+y=21\) .... (1) That was the first equation. The second equation can be found by considering the total number of words. \(1800y+2400x=48000\) ... (2)

OpenStudy (anonymous):

Solve the two equations. the value of x is the number of pages in smaller type, and that's your answer.

OpenStudy (anonymous):

You will find that the number of the smaller pages x=17 pages.

OpenStudy (anonymous):

i have forgotten how to solve my x and y...please help it has been a long time for me

OpenStudy (anonymous):

From equation (1), we have \(y=21-x\). Substitute that in (2) and get x. \(1800(21-x)+2400x=48000 \implies 37800+600x=48000\) \(\implies 600x=10,200 \implies x=17\).

OpenStudy (anonymous):

got it thank you so much